There are three equally likely states of nature (High, Medium, and Low demand). If the large factory will post profits of $60,000, $25,000, and -$10,000 under these states of nature, respectively, what is the EMV of the factory?
1
Expert's answer
2020-12-07T21:01:42-0500

If they are equally likely, EMV is

"\frac 13 \cdot 60000+\frac 13 \cdot25000-\frac 13 \cdot10000=\frac 13 \cdot 75000=25000."